log6(x-1)+log6(x+4)=2
Answer by nerdybill(7384)

log6(x-1)+log6(x+4)=2
log6(x-1)(x+4)=2
(x-1)(x+4)=6^2
x^2+3x-4=36
x^2+3x-40=0
(x+8)(x-5) = 0
x = {-8,5}
throw out the -8 (extraneous) leaving:
x = 5